# references.bib

### v1

@@ 1 Introduction
Residual connections play an important role in modern neural network architectures because they stabilize the training of deep neural networks and improve model convergence and quality. Since their usage in the ResNet architecture [he2016deep], residual connections have been widely adopted in both convolutional neural networks and transformer architectures across various domains, including natural language processing [vaswani2017attention], audio recognition [gong21b_interspeech], and computer vision [dosovitskiy2020image].
@@
In order to confirm the existence and significance of the dilution phenomenon we ask a simple question: Can residual networks easily learn to recover the input? This should be a basic task expected of any generative model — otherwise there would be information loss. However, if our dilution hypothesis is true, the answer would be negative. To test this, we create a neural network consisting of a number of low-rank layers, and add residual connections in order to mitigate the bottlenecks introduced by the low ranks. The resulting model is full-rank. We compare this model with another model that employs learnable residual connections, as in DenseFormer [pagliardini2024denseformer], which we later also call GRN-v1, since it is the starting point of our generalizations. In Figure 1 we see the results of the two models on two tasks: learning the identity transformation and learning a random linear transformation. Perhaps surprisingly, we observe that the residual network is unable to fully reconstruct the input even after seeing $10^{3}$ batches ($10^{5}$ examples), while the model with learnable residual weights is able to reach extremely small loss values, even with 100x fewer examples. This confirms that ResNet does not address neural network bottlenecks in a satisfactory way, even though it learns a full-rank transformation, and underscores the importance of using learnable residual weights to increase model capacity.
@@
[…]
We additionally provide empirical results to support the theoretical findings and demonstrate the effectiveness of DCA. Experiments on language modeling tasks demonstrate that DCA consistently outperforms the standard transformer architectures in terms of both perplexity and training efficiency. DCA achieves lower perplexity for a given parameter budget and training time. Furthermore, DCA exhibits improved training stability, mitigating the occurrence of loss spikes frequently observed while training large models.
@@ 2 Related work
[…]
@@ 3 Method
Notation. We denote a residual function by $f_{t}:\mathbb{R}^{d}\to\mathbb{R}^{d}$, where $t$ is the layer index and $d$ the feature dimension. As an example, in a multi-layer perceptron residual network (MLP-ResNet), we have $f_{t}(\bm{x})={\bm{V}}_{t}\sigma(\bm{W}_{t}\bm{x})$ with $\bm{W}_{t}\in\mathbb{R}^{k\times d}$, ${\bm{V}}_{t}\in\mathbb{R}^{d\times k}$ and $\sigma$ is a nonlinear function, such as sigmoid or ReLU, that is applied component-wise. Then, the $t$-th residual block outputs $g_{t+1}(\bm{x})$, defined recursively as
@@ 3.1 Generalized Residual Networks (GRN)
We propose three generalizations of ResNets by considering weighted linear combinations of previous layer outputs. The parameters of the modules and the generalizations are all optimized during training using the AdamW optimizer [loshchilov2017decoupled].
@@
with $\bm{b}_{t}\in\mathbb{R}^{t\times 1}$ which is initialized as all ones and optimized with the rest of the model parameters during training. This setting has been previously explored in the DenseFormer paper [pagliardini2024denseformer].
@@ 4.1 Low-rank linear model
For the low rank linear model we have: $\bullet\quad\mathcal{C}_{{\rm base}}=\{\bm{x}\mapsto{\bm{M}}\bm{x}:\;\operatorname{rank}({\bm{M}})\leq\min(r_{t})_{t=1}^{T}\}$.
@@ 4.2 Trade-off between test error and model complexity
where $\{\Delta_{ii}\}_{i=1}^{d}$ are the diagonal entries of $\bm{\Delta}$.
@@
$G_{2}$ $\geq(d-r_{*})\lambda_{\min}^{2}-(\sqrt{1+r_{*}}-1)^{2}(\lambda_{\max}^{2}-\lambda_{\min}^{2})\,.$
@@
Consider a ResNet with collective rank $r_{*}=\sum_{t=1}^{T}r_{t}<d$. A GRN-v1 or GRN-v2 model can achieve a smaller test error with collective rank $r^{\prime}_{*}$, where $r^{\prime}_{*}:=\frac{r_{*}-d\kappa^{2}}{1-\kappa^{2}}<r_{*}$.
@@ 4.3 Insights from the analysis
Role of target task complexity. Note that $\kappa=\lambda_{\min}/\lambda_{\max}\in[0,1]$ is a measure of complexity of the target task. Specifically, as $\kappa$ decreases, the matrix ${\bm{A}}$ becomes closer to a low rank matrix, and hence learning it with low rank models becomes easier. Observe that the thresholds given by the right hand side of (4.1) and (4.2) are increasing in $\kappa$, i.e., for more complex tasks we see a wider range of collective rank where GRNs outperforms the trade-off achieved by ResNet. Another way to interpret Theorem 4.3(i) is that for a fixed target task (and so fixed $\kappa$), if the collective rank $r_{*}$ is above this threshold, the ResNet is already rich enough that it is hard to improve upon its trade-off.
Role of collective rank. Observe that the lower bound on the gains $G_{1}$, $G_{2}$ given by Theorem 4.3(ii) are decreasing in $r_{*}$. In other words, when the collective rank $r_{*}$ of ResNet becomes smaller, the level of information dilution occurring in ResNet increases, giving GRNs a better leverage to improve model perplexity with the same number of parameters.
Role of input dimension. Note that the upper bounds on $r_{*}$ given by (4.1) and (4.2) increase with the input dimension $d$. Furthermore, the lower bounds on the gains $G_{1}$, $G_{2}$, given in Theorem 4.3(ii) also increase with $d$. Therefore, for larger input dimensions, we have both a wider range for $r_{*}$ where GRNs outperforms the trade-off achieved by ResNet, and moreover, we obtain a larger gain in reducing model error.
@@ 4.4 Extension to nonlinear models
We recall the definition of Bottleneck rank from [jacot2022implicit]. For a function $f:\Omega\mapsto\mathbb{R}^{d}$, its Bottleneck rank, denoted by $\operatorname{rank}_{BN}(f,\Omega)$ is the smallest integer $k$ such that $f$ can be factorized as $f=h\circ g$ with inner dimension $k$ (i,e, $g:\Omega\mapsto\mathbb{R}^{k}$ and $h:\mathbb{R}^{k}\mapsto\mathbb{R}^{d}$) It is also closely related to the Jacobian rank of a function defined as $\operatorname{rank}_{\bm{J}}(f)=\max_{\bm{x}\in\Omega}\operatorname{rank}[\bm{J}f(\bm{x})]$. In general, $\operatorname{rank}_{\bm{J}}(f)\leq\operatorname{rank}_{BN}(f)$, but for functions of the form $f=\psi\circ{\bm{A}}\circ\phi$ (for a linear map ${\bm{A}}$ and two bijections $\psi$ and $\phi$), we have $\operatorname{rank}_{\bm{J}}(f)=\operatorname{rank}_{BN}(f)=\operatorname{rank}({\bm{A}})$. These two notions of rank satisfy the following properties [jacot2022implicit]:
@@
$\bullet\quad\mathcal{C}_{{\rm GRN-v2{}}}\subseteq\left\{{\bm{D}}+f:\;\operatorname{rank}_{BN}(f)\leq r_{*},\,{\bm{D}}\text{ is diagonal}\right\}$.
@@ 5 Experiments
[…]
@@
Model width scaling. Our theoretical results indicate that the benefit of GRN is inversely related to the rank of the model. With this experiment, we validate whether the theoretical results carry over to the transformer architecture by varying the model width. Table 2 shows the final perplexity of a 24-layer model with an embedding dimension ranging from 64 till 1024, pre-trained on LM1B. The delta column, with the difference between the transformer and DCA, shows that the benefit of DCA is reduced as the width of the model increases, which is consistent with our theoretical results. These results are in contrast with the depth scaling results, where the improvement of DCA is maintained for deeper models.
@@
Retrofitting pre-trained models. Since our method is identical to a standard residual network at initialization, adding DCA to a pre-trained model does not alter its function. In Table 4, we compare continuing training the pre-trained model with adding DCA to the pre-trained model. Incorporating DCA results in a perplexity improvement of 0.23 after 60k extra training steps, compared to just 0.015 for the transformer. Thus, pre-trained models with a residual architecture can also benefit from incorporating DCA.
@@
500k 18.963 18.982 0.019
500k + 20k 18.978 18.803 -0.175
500k + 40k 18.950 18.775 -0.175
500k + 60k 18.948 18.752 -0.196
[…]
Table 5: Perplexity (PPL) and parameter count on LM1B with a 6-layer model, comparing DCA with related work.
@@
Transformer 49.65M 18.961
LAuReL-PA 49.75M 18.954
1x1-DenseFormer 49.65M 18.669
Hyper-Connections (dynamic) 49.68M 18.633
DCA (ours) 49.73M 17.998
@@
Transformer 49.65M 18.961
GRN-v1 49.65M 18.669
GRN-v2 49.66M 18.378
GRN-v3 49.68M 18.332
DCA 49.73M 17.998

### v2

# DeepCrossAttention: Supercharging Transformer Residual Connections

@@ Abstract
Residual Network, Cross Attention, ResNet, Transformer
@@ 1 Introduction
Residual connections play an important role in modern neural network architectures because they stabilize the training of deep neural networks and improve model convergence and quality. Since their usage in the ResNet architecture (He et al., 2016), residual connections have been widely adopted in both convolutional neural networks and transformer architectures across various domains, including natural language processing (Vaswani, 2017), audio recognition (Gong et al., 2021), and computer vision (Dosovitskiy et al., 2021).
@@
In order to confirm the existence and significance of the dilution phenomenon we ask a simple question: Can residual networks easily learn to recover the input? This should be a basic task expected of any generative model — otherwise there would be information loss. However, if our dilution hypothesis is true, the answer would be negative. To test this, we create a neural network consisting of a number of low-rank layers, and add residual connections in order to mitigate the bottlenecks introduced by the low ranks. The resulting model is full-rank. We compare this model with another model that employs learnable residual connections, as in DenseFormer (Pagliardini et al., 2024), which we later also call GRN-v1, since it is the starting point of our generalizations. In Figure 1 we see the results of the two models on two tasks: learning the identity transformation and learning a random linear transformation. Perhaps surprisingly, we observe that the residual network is unable to fully reconstruct the input even after seeing $10^{3}$ batches ($10^{5}$ examples), while the model with learnable residual weights is able to reach extremely small loss values, even with 100x fewer examples. This confirms that ResNet does not address neural network bottlenecks in a satisfactory way, even though it learns a full-rank transformation, and underscores the importance of using learnable residual weights to increase model capacity.
@@
We analyse our generalization of the residual network theoretically by focusing on a linear low-rank model. We show that DCA achieves a better trade-off between accuracy and model size when the ratio of the collective ranks of the layers to the ambient dimension is below a threshold, which depends on the complexity of the target task. In addition, the improvement in this trade-off can itself be characterized as a function of the collective ranks of the layers, ambient dimension and the complexity of the target task. We extend this insight to nonlinear models by working with the notion of bottleneck rank, proposed by Jacot (2023).
We additionally provide empirical results to support the theoretical findings and demonstrate the effectiveness of DCA. Experiments on language modeling and image classification tasks demonstrate that DCA consistently outperforms the standard transformer architectures in terms of perplexity, accuracy and training efficiency. DCA achieves lower perplexity for a given parameter budget and training time. Furthermore, DCA exhibits improved training stability, mitigating the occurrence of loss spikes frequently observed while training large models.
@@ 2 Related Work
Highway networks enable each layer to interpolate dynamically between its output $f(\bm{x})$ and its input $\bm{x}$ using a gating mechanism (Srivastava et al., 2015). Residual connections (He et al., 2016) popularized the direct flow of information from earlier to later layers using identity shortcuts. These innovations proved crucial in stabilizing training and allowing for the construction of significantly deeper networks. Building upon this concept, DenseNet (Huang et al., 2017) further enhanced information flow by concatenating the outputs of all preceding layers to each layer’s input.
The following methods are the ones most similar to ours. They all build on the idea of DenseNet but apply an efficient aggregation of the previous layer outputs instead of concatenating them. DenseFormer (Pagliardini et al., 2024) performs the aggregation as a learned linear combination of the previous layer outputs. To reduce the computational load, they propose to apply their method only on a subset of the possible layer connections. Building on DenseFormer, LAuReL (Menghani et al., 2024) presents three aggregation functions, the best performing one applies a learned low-rank transformation to the previous layer outputs before the learned linear combination. Zhu et al. (2024) take a different approach with Hyper-Connections, they consider a fixed-size stack where layer outputs are added into with a learned weight for every slot of the stack. Before each layer, the stack is mixed by a matrix multiplication with a learned weight matrix. The input to a layer is then obtained by a learned linear combination of the stack, instead of accessing the previous layer outputs directly. They also present a dynamic version of their method where the weights are derived from the inputs.
@@ 3 Method
Notation. We denote a residual function by $f_{t}:\mathbb{R}^{d}\to\mathbb{R}^{d}$, where $t$ is the layer index and $d$ the feature dimension. As an example, in a multi-layer perceptron residual network (MLP-ResNet), we have $f_{t}(\bm{x})={\bm{V}}_{t}\sigma(\bm{W}_{t}\bm{x})$ with $\bm{W}_{t}\in\mathbb{R}^{k\times d}$, ${\bm{V}}_{t}\in\mathbb{R}^{d\times k}$ and $\sigma$ is a nonlinear function, such as sigmoid or ReLU, that is applied component-wise. Then, the $t$-th residual block outputs $g_{t+1}(\bm{x})$, are defined recursively as
@@ 3.1 Generalized Residual Networks (GRN)
We propose three generalizations of ResNets by considering weighted linear combinations of previous layer outputs. The parameters of the modules and the generalizations are all optimized during training using the AdamW optimizer (Loshchilov & Hutter, 2017).
@@
with $\bm{b}_{t}\in\mathbb{R}^{t\times 1}$ which is initialized as all ones and optimized with the rest of the model parameters during training. This setting has been previously explored in the DenseFormer paper (Pagliardini et al., 2024).
@@ 4.1 Low-rank linear model
GRN-v3. In this case, we have $\widehat{\bm{y}}=({\bm{G}}_{T+1}\odot(\bm{b}_{T+1}+\bar{\bm{w}}_{T+1}))\bm{1}$, where $\bm{b}_{T+1}$ is $d\times(T+1)$ matrix and $\bar{\bm{w}}_{T+1}$ is $d$ dimensional vector as described in Section 3. We denote by $\mathcal{C}_{{\rm GRN-v3{}}}$ the class of functions that can be represented by such architecture.
@@
For the low rank linear model we have:
@@
$\bullet\quad\mathcal{C}_{{\rm GRN-v3{}}}\supset\{\bm{x}\mapsto({\bm{D}}+{\bm{M}})\bm{x}+\sigma(\bm{w}^{\sf T}\bm{x})\bm{x}:\;\operatorname{rank}({\bm{M}})\leq r_{*},{\bm{D}}\text{ is diagonal},\bm{w}\in\mathbb{R}^{d\times 1}\}$.
@@ 4.2 Trade-off between test error and model complexity
where $\{\Delta_{ii}\}_{i=1}^{d}$ are the diagonal entries of $\bm{\Delta}$. In addition,
${\sf Err}^{*}(\mathcal{C}_{\rm GRN-v3{}})\leq\left\|\bm{\Delta}\right\|_{F}^{2}-\max\{\mathcal{T}_{1},\mathcal{T}_{2}\}\,,$
$\mathcal{T}_{1}:=\frac{1}{d-r_{*}}{{\operatorname{trace}}}(\bm{\Delta})^{2}+\frac{\nu_{\max}^{2}}{\pi(d+1)},$
$\mathcal{T}_{2}:=\sum_{i=1}^{d}\Delta_{ii}^{2}+\frac{\tilde{\nu}_{\max}^{2}}{\pi(d+1)}\,.$
Here $\nu_{\max}$ denotes the maximum eigenvalue of $\frac{\bm{\Delta}+\bm{\Delta}^{\sf T}}{2}-\frac{1}{d-r_{*}}{{\operatorname{trace}}}(\bm{\Delta}){\bm{U}}_{r*,\perp}{\bm{U}}_{r_{*},\perp}^{\sf T}$ and $\tilde{\nu}_{\max}$ denotes the maximum eigenvalue of $\frac{\bm{\Delta}+\bm{\Delta}^{\sf T}}{2}-\text{diag}(\bm{\Delta})$.
@@
Also, GRN-v3 achieves a better trade-off compared to ResNet, if
$r_{*}\leq(1+\eta(\sqrt{\eta^{2}+d}-\eta))^{2}-1.6\,,$ (4.3)
where $\eta=\sqrt{\frac{(\kappa(1+\xi_{0})-\xi_{0})^{2}+\xi_{0}}{1+\xi_{0}}}$ and $\xi_{0}=\frac{1}{\pi(d^{2}-1)}$.
@@
$G_{2}$ $\geq(d-r_{*})\lambda_{\min}^{2}-(\sqrt{1+r_{*}}-1)^{2}(\lambda_{\max}^{2}-\lambda_{\min}^{2})\,,$
$G_{3}$ $\geq\left(d-\frac{1}{\pi(d+1)}-r_{*}\right)\lambda_{\min}^{2}+\frac{1}{2\pi(d+1)}\lambda_{\max}^{2}$
$\quad-(\sqrt{1.6+r_{*}}-1)^{2}(\lambda_{\max}^{2}-\lambda_{\min}^{2})\,.$
@@
Consider a ResNet with collective rank $r_{*}=\sum_{t=1}^{T}r_{t}<d$. A GRN-v1 or GRN-v2 model can achieve a smaller test error with collective rank $r^{\prime}_{*}$, where $r^{\prime}_{*}:=\frac{r_{*}-d\kappa^{2}}{1-\kappa^{2}}<r_{*}$. Likewise, a GRN-v3 model achieve a smaller test error with collective rank $\tilde{r}_{*}$, where $\tilde{r}^{\prime}_{*}:=\frac{r_{*}-d\eta^{2}}{1-\eta^{2}}<r_{*}$, with $\eta=\sqrt{\frac{(\kappa(1+\xi_{0})-\xi_{0})^{2}+\xi_{0}}{1+\xi_{0}}}$ and $\xi_{0}=\frac{1}{\pi(d^{2}-1)}$.
@@ 4.3 Insights from the analysis
Role of target task complexity. Note that $\kappa=\lambda_{\min}/\lambda_{\max}\in[0,1]$ is a measure of complexity of the target task. Specifically, as $\kappa$ decreases, the matrix ${\bm{A}}$ becomes closer to a low rank matrix, and hence learning it with low rank models becomes easier. Observe that the thresholds given by the right hand side of (4.1)-(4.3) are increasing in $\kappa$, i.e., for more complex tasks we see a wider range of collective rank where GRNs outperforms the trade-off achieved by ResNet. Another way to interpret Theorem 4.3(i) is that for a fixed target task (and so fixed $\kappa$), if the collective rank $r_{*}$ is above this threshold, the ResNet is already rich enough that it is hard to improve upon its trade-off.
Role of collective rank. Observe that the lower bound on the gains $G_{1}$, $G_{2}$, $G_{3}$ given by Theorem 4.3(ii) are decreasing in $r_{*}$. In other words, when the collective rank $r_{*}$ of ResNet becomes smaller, the level of information dilution occurring in ResNet increases, giving GRNs a better leverage to improve model perplexity with the same number of parameters.
Role of input dimension. Note that the upper bounds on $r_{*}$ given by (4.1) to (4.3) increase with the input dimension $d$. Furthermore, the lower bounds on the gains $G_{1}$, $G_{2}$, $G_{3}$ given in Theorem 4.3(ii) also increase with $d$. Therefore, for larger input dimensions, we have both a wider range for $r_{*}$ where GRNs outperforms the trade-off achieved by ResNet, and moreover, we obtain a larger gain in reducing model error.
@@ 4.4 Extension to nonlinear models
We recall the definition of Bottleneck rank from (Jacot, 2023). For a function $f:\Omega\mapsto\mathbb{R}^{d}$, its Bottleneck rank, denoted by $\operatorname{rank}_{BN}(f,\Omega)$ is the smallest integer $k$ such that $f$ can be factorized as $f=h\circ g$ with inner dimension $k$ (i,e, $g:\Omega\mapsto\mathbb{R}^{k}$ and $h:\mathbb{R}^{k}\mapsto\mathbb{R}^{d}$) It is also closely related to the Jacobian rank of a function defined as $\operatorname{rank}_{\bm{J}}(f)=\max_{\bm{x}\in\Omega}\operatorname{rank}[\bm{J}f(\bm{x})]$. In general, $\operatorname{rank}_{\bm{J}}(f)\leq\operatorname{rank}_{BN}(f)$, but for functions of the form $f=\psi\circ{\bm{A}}\circ\phi$ (for a linear map ${\bm{A}}$ and two bijections $\psi$ and $\phi$), we have $\operatorname{rank}_{\bm{J}}(f)=\operatorname{rank}_{BN}(f)=\operatorname{rank}({\bm{A}})$. These two notions of rank satisfy the following properties (Jacot, 2023):
@@
$\bullet\quad\mathcal{C}_{{\rm GRN-v2{}}}\subseteq\{g:\,g(\bm{x})={\bm{D}}\bm{x}+f(\bm{x}):\;\operatorname{rank}_{BN}(f)\leq r_{*},\,{\bm{D}}\text{ is diagonal}\}$.
@@ 5 Experiments
We conduct experiments on language modeling and image classification tasks to evaluate the effectiveness of DCA and to validate our theoretical insights. For the language modeling tasks, the performance of DCA is compared against the standard transformer (Vaswani, 2017) on the LM1B (Chelba et al., 2013) and C4 (Raffel et al., 2020a) datasets. Unless stated otherwise, each model has an embedding dimension of 512 and an MLP dimension of four times the embedding dimension. By default, DCA uses a stack of all the previous layer outputs as input to the GRNs. When DCA includes only the first and last-$k$ layer outputs explicitly in the input stack (see Section 3.1), then this is denoted as $k$-DCA.
Each model is trained with a sequence length of 128 and a batch size of 2048 over 64 TPUs for 500k steps, totaling 131B tokens. We use the AdamW optimizer (Loshchilov & Hutter, 2017) with $\beta_{1}=0.9$, $\beta_{2}=0.98$, a weight decay of $0.1$, and a learning rate of $0.0016$ with 1000 warmup steps and an inverse square root schedule (Raffel et al., 2020b).
Model depth scaling. For the first experiment, we pre-train a transformer and DCA on LM1B. We increase the model depth from 6 to 42 layers and show the relation between perplexity (Jelinek et al., 1977) and model size in Figure 6. The figure shows that DCA obtains a lower perplexity for a given parameter budget. Notably, the 30-layer DCA model obtains a better perplexity than the 42-layer transformer, making DCA more parameter-efficient than adding layers.
@@
Model width scaling. Our theoretical results indicate that the benefit of GRN is inversely related to the rank of the model. With this experiment, we validate whether the theoretical results carry over to the transformer architecture by varying the model width. Table 2 shows the final perplexity of a 12-layer model with an embedding dimension ranging from 64 till 1024, pre-trained on LM1B. The delta column, with the difference between the transformer and DCA, shows that the benefit of DCA is reduced as the width of the model increases, which is consistent with our theoretical results. These results are in contrast with the depth scaling results, where the improvement of DCA is maintained for deeper models.
@@
Retrofitting pre-trained models. Since our method is identical to a standard residual network at initialization, adding DCA to a pre-trained model does not alter its function. In Table 4, we compare continuing training the pre-trained model with adding DCA to the pre-trained model. Incorporating DCA results in a perplexity improvement of 0.19 after 60k extra training steps, compared to just 0.02 for the transformer. Thus, pre-trained models with a residual architecture can also benefit from incorporating DCA.
@@
500k 18.98$\pm$0.01 18.98$\pm$0.01 0.00
500k + 20k 18.96$\pm$0.02 18.81$\pm$0.01 -0.15
500k + 40k 18.96$\pm$0.01 18.79$\pm$0.03 -0.17
500k + 60k 18.96$\pm$0.01 18.79$\pm$0.04 -0.17
Training stability. The occurrence of loss spikes is a problem when training large models as they can disrupt an expensive training run (Chowdhery et al., 2023). In Figures 7 and 8, we indeed observe clear loss spikes with the transformer model. Interestingly, training DCA is more stable, showing no significant loss spikes even for large models. This constitutes an important benefit of DCA.
Comparison with related work. We compare the perplexity of DCA with those obtained by the recent related works LAuReL (Menghani et al., 2024), DenseFormer (Pagliardini et al., 2024), and Hyper-Connections (dynamic) (Zhu et al., 2024) in Table 5. DCA improves upon the prior best method, hyper-connections, with a difference in perplexity of 0.59, which is the biggest improvement among the methods.
Table 5: Perplexity (PPL) and parameter count on LM1B using a 6-layer model, comparing DCA with related work.
@@
Transformer 49.65M 18.98$\pm$0.01
LAuReL-PA 49.75M 18.99$\pm$0.05
1x1-DenseFormer 49.65M 18.80$\pm$0.11
Hyper-Connections 49.68M 18.65$\pm$0.03
DCA (ours) 49.73M 18.06$\pm$0.01
Table 6: Perplexity (PPL) on C4 using a 13-layer model and embedding dimension $1$, comparing DCA with related work. The baseline model has roughly 179M parameters.
Method PPL
Transformer 21.534
LAuReL-PA 20.951
1x1-DenseFormer 21.168
Hyper-Connections (stack size=4) 21.077
Hyper-Connections (stack size=10) 20.718
$8$-DCA (ours) 20.392
@@
Transformer 49.65M 18.98$\pm$0.02
GRN-v1 49.65M 18.80$\pm$0.11
GRN-v2 49.66M 18.43$\pm$0.04
GRN-v3 49.68M 18.41$\pm$0.10
DCA 49.73M 18.06$\pm$0.01
ImageNet classification. In addition to the language modelling experiments, we also experiment with image classification using the ImageNet dataset and the vision transformer (ViT) model (Dosovitskiy et al., 2021). Since the ViT model is transformer-based, DCA can be incorporate in the same way as for the language models presented earlier. In Table 8, we present the results on the ViT-S/16 model (22M parameters) and follow the experimental setup by Beyer et al. (2022). The results show a 0.7% improvement in classification accuracy, demonstrating that DCA effectively generalizes to the vision domain.
Table 8: Loss and Accuracy on ImageNet classification.
Method Loss Accuracy
ViT 0.5698 76.4
ViT + DCA (ours) 0.5284 77.1